OpenStudy (anonymous):

Please help me it say to Simplify this radical and to remember that i = sqrt(-1) problem => √(-17) *√(-48)

OpenStudy (anonymous):

Hint:\[\sqrt{-17}\sqrt{-48}=i\sqrt{17}\ i\sqrt{48}=\cdots\]

OpenStudy (anonymous):

thank you